OpenStudy (anonymous):

Can someone please check my answers and help me on a few unanswered questions?

OpenStudy (anonymous):

For the function ƒ(x) = 10x2 + x, evaluate and simplify the expression f(a+h)-f(a) / h

OpenStudy (anonymous):

I got f(a+h)-f(a) / h = 20a+1

OpenStudy (anonymous):

2. Find ƒ−1 for the function f(x) = 23-3x / 6x I got f^-1 (x) = 23/6x-3

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):



OpenStudy (anonymous):

for both of them?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u know any math websites that really help? like wolfgram for example.

OpenStudy (anonymous):

ya u r

OpenStudy (anonymous):

yes go to http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

:0 thatnks

OpenStudy (anonymous):

You are correct on both!

OpenStudy (anonymous):

Yay

OpenStudy (anonymous):

Any time!